OpenStudy (anonymous):

Write the equation of a line that has a slope of -2/3 and passes through the point (9, 3)

OpenStudy (anonymous):

\[ y = mx +b \\ m = -\frac{2}{3} \\ y = -\frac{2}{3}x + b \\ \text{For point (9,3):} \quad x = 9 \quad \quad y=3 \\ 3 = -\frac{2}{3} \cdot 9 +b \\ 3 = -6 +b \\ b = 9 \] So \[ y = -\frac{2}{3}x + 9 \]

OpenStudy (anonymous):

Thank you so much :))))